your retirement pension is an annuity with a guarnteeed return of 4% interest compounded monthly. you would like to retire with a pension of 3000 per month for 25 years.
A) how much money needs to be in account at the time of retirement in order to fund these withdraws
B) how much must be deposited each month into the accoint during the 45 years leading up to retirement in order to achieve this
Answer #1

A) PMT = 3,000

I/Y = 4/12 = 0.3333333333

FV = 0

N = 25*12 = 300

CPT PV

PV = -568,357.4489

The amount that needs to be in the account at the time of the retirement in order to fund these withdrawals is $568,357.4489

B) N = 45*12 = 540

I/Y = 0.3333333333

PV = 0

FV = 568,357.4489

CPT PMT

PMT = -376.5278831

The amount that needs to be deposited each month into the accoint during the 45 years leading up to retirement in order to achieve $568,357.4489 is $376.5278831

Can you please upvote? Thank You :-)